2. (25pts.) Find Ref and Im f: f(2)=4z2 + 5z +2i-1
Answer #1

Here 2 C aNaw, ut wherre o TRa Mai. + - - (4-445-1)(ye52) . Ref= 4-4j45x- Im=yf5yta